q1

ABC has the following capital structure. What is the WACC for the company?

 

Debt:

Bond 1. 400,000 bonds with a coupon rate of 6% (paid semi-annually), a price quote of 110.0 and have 20 years to maturity.

 

 

Bond 2. 300,000 zero coupon bonds (semi-annual compounding) with a price quote of 40.0 and 25 years until maturity.

 

Preferred stock:

600,000 shares of 4 percent preferred stock with a current price of $60, and a par value of $100.

 

Common stock:

12,000,000 shares of common stock, with a price of $80, and a beta of 1.3.

 

Market:

The corporate tax rate is 40 percent, the market risk premium is 9%, and the risk-free rate is 4%.

 

 

 

 

Question 2

The ABC stock has a beta of 1.4. The company just paid a dividend of $2.0, and the dividend is expected to grow at 6% per year, indefinitely. The expected market return is 14%, and the risk-free is 4%. The most recent stock price for ABC is $60. Calculate the cost of equity using the SML method.

 



1
Expert's answer
2022-02-03T13:56:32-0500

Q1

"B1=400 000\times110=44 000 000"

"B2=300000\times40=12 000 0000"

"Sp=600 000\times60=36 000 000"

"Sc=12 000 000\times80=960 000 000"


capital=44+12+36+960=1052


let's N=120

"YTM1=\frac{7.2+\frac{120-110}{20}}{\frac{120+110}{2}}=\frac{7.7}{115}=0.0669"

"YTM1=0.0669\times2=0.1338"


let's N=50

"YTM2=\frac{50}{40}^{\frac{1}{25}}-1=0.0089"

"YTM2=0.0089\times2=0.0178"


"CARM=4+1.3(9-4)=10.5"


"r=\frac{4}{60}=0.0667"


"WACC=\frac{44}{1052}\times13.38(1-0.4)+\frac{12}{1052}\times1.78(1-0.4)+\frac{36}{1052}\times6.67+\frac{960}{1052}\times10.5=10.16"


Q2

"CARM=4+1.4(14-4)=18"